OpenStudy (anonymous):

if f(x)=x=4 and g(x)=2x-3, Find g-f(2)

OpenStudy (unklerhaukus):

f(x) = x-4, g(x) = 2x-3 (g-f)(x) = g(x)-f(x) = (x-4)-(2x-3) (g-f)(2) = g(2)-f(2) = ((2)-4)-(2(2)-3)